OpenStudy (anonymous):

integral of 3x^2-2x+3 on the bound 2,1... need to know how

OpenStudy (anonymous):

\[ \int\limits_{1}^{2} (3-2 x+3 x^2) dx\] Integrate the sum term by term and factor out constants: \[3 \int\limits_{1}^{2} x^2 dx+ 3\int\limits_{1}^{2} dx-2 \int\limits_{1}^{2} x dx\] The integral of x^2 is x^3/3: \[x^3+ 3\int\limits_{1}^{2} dx-2 \int\limits_{1}^{2} x dx\] The integral of x is x^2/2: \[x^3-x^2+ 3\int\limits_{1}^{2} dx\] The integral of dx is x: \[[x^3-x^2+3x]\] Now taking it from 1 to 2: \[[(2)^3-(2)^2+3(2)]-[(1)^3-(1)^2+3(1)]=7\] So your final answer is 7.

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yes it does...

OpenStudy (anonymous):

but instead for going term by term couldn't you take the derivative of the function and then evaluate it for the two bound numbers

OpenStudy (anonymous):

you can take the integral at the same time like this integral of 3x^2-2x+3 on the bound 2,1. =(3x^3)/3 -(2x^2)/2+3x] =x^3 -x^2 +3x] now taking x from 2 to 1 =[2^3 -2^2 +3(2)]-[1^3 -1^2 +3(1] =8-4+6-(3)=7 ans.......